With Picos, we lose the ability to dither the input beam as well as align the beam with the IFO locked. And the active TT will still have hysteresis, but also actuators. Once in vacuum, I'm not sure how we adjust them - what's the error signal for PR2/PR3 ?
If the interferometer is aligned, why not just pump down now? I'm not sure that we have evidence of TT hysteresis issues once people stop touching them.